# ReactionMamba: Generating Short & Long Human Reaction Sequences

### v1

@@ Abstract
We present ReactionMamba, a novel framework for generating long 3D human reaction motions. ReactionMamba integrates a motion VAE for efficient motion encoding with Mamba-based state-space models to decode temporally consistent reactions. This design enables ReactionMamba to generate both short sequences of simple motions and long sequences of complex motions, such as dance and martial arts. We evaluate ReactionMamba on three datasets—NTU120-AS, Lindy Hop, and InterX—and demonstrate competitive performance in terms of realism, diversity, and long-sequence generation compared to previous methods, including InterFormer, ReMoS, and Ready-to-React, while achieving substantial improvements in inference speed.
@@ I Introduction
In recent years, there has been impressive progress in human motion generation under various user-specified conditions, such as music pieces [64] and text prompts [40]. However, most of these approaches focus on a single-person setting [7, 39] and, as a result, neglect one essential aspect of human motion: rich human-to-human interactions . One important aspect of human interaction is that the action of one person will usually elicit a reaction from another person. Synthesizing human-to-human interactions introduces advanced capabilities for character animation tools and software, with applications across commercial and entertainment media [24], interactive mixed and augmented reality [18], social robotics [59], and avatar interaction [8].
@@
This led recent research to take interest in the task and propose various architectures to generate reaction motions [20, 9, 43]. However, human interaction reactions present a more challenging task due to the motion reactions, which are highly complex and have low periodicity. They require extremely accurate modeling of both spatial and temporal features, without which the generated motion looks unrealistic. While current state-of-the-art methods can now generate reactive motion for simple and short actions they still struggle with longer sequences. Indeed short sequences often include simple motion such as boxing or kicking that can be more easily modeled due to a direct body contact leading to a clear reaction. Longer sequences however often include more complex interactions, such as dancing, where the relation between the action and the reaction is not always as direct and obvious. Modeling longer interaction sequences requires specific architectures, different from those focused on short sequences, to achieve high quality reaction generation. Precursor works on reaction generation have represented the task as a translation problem which directed them towards the Transformer [9]. However, this led to low diversity generation and limited performance in the long term. The self-attention mechanism computes interactions between every pair of tokens in the input sequence. This leads to quadratic time complexity in relation to the sequence length, which becomes a bottleneck for processing long sequences. As a result, it is essential to explore a new architectural paradigm that can handle long-range dependencies while maintaining linear computational complexity to support motion generation tasks. Recent advances have led to renewed interest in state-space models (SSMs) [23]. SSMs are a class of mathematical models used for modeling dynamic systems over time, often in fields like control theory, signal processing, and time series analysis. They have several advantages, particularly in their ability to model temporal dependencies for long sequences [23]. The efficiency in managing long sequences primarily due to its implementation of convolutional computations, which efficiently process sequential data. Convolutional operations reduce the complexity associated with handling long-range dependencies by applying local filters to the sequence such as Mamba [22]. In this paper, we propose ReactionMamba (Figure 2) a new architecture for reaction motion generation that matches the state-of-the-art while increasing inference speed by a hundredfold. The Figure LABEL:fig:teaser shows some examples results obtained by our framework.
@@
  • We introduce a simple yet effective framework, named ReactionMamba, which is a pioneering method integrates Mamba into human motion reaction generation.
  • ReactionMamba is VAE [29] made up of three interconnected modules, an encoder, a conditioning module, and a decoder. The SSM based encoder encodes the reaction into a sequence of continuous latent variables in an efficient way, and a SSM based decoder is then designed to capture information hidden in the latent representations of the pose, combined with the action motion sequence and the initial frame of the reaction and to reconstruct the reaction.
  • Unlike conventional deterministic methods, our method uses random latent sampling at inference, conditioned by the projected action sequence and the projected initial reaction pose, providing diverse and realistic reaction sequences.
  • The ReactionMamba framework demonstrated exceptional performance on the action-reaction generation task for simple action-reaction categories [55] and  [54], two-person interactions such as LindyHop dancing and Ninjutsu martial arts [20], and acrobats performing figures [52]. ReactionMamba achieved state-of-the-art generation quality, diversity while greatly improving processing speed.
@@ II Related work
Single-Person Human Motion Synthesis. Single-human motion synthesis has seen remarkable progress in recent years. To cite some examples, conditional VAEs [39, 11] have been proposed to map textual descriptions to human motion. More recent approaches, such as EDGE [47] and Dancing to Music [30], introduce frameworks for the generation of music-conditioned motion. However, these methods often require additional conditions to effectively model emotional expressiveness. In addition, these approaches focus on the generation of single-human while ReactMamba learns the synchronization between two interacting individuals directly from one person’s motion, without the need for additional prompts or labels.
@@
Reactive Motion Generation. Reactive motion generation focuses on generating human motion in response to external agents. Reactive motion generation is inherently more challenging than generating a single human motion sequence. In [58], a motion graph synthesis technique is proposed that generates gestures and body motions in a dyadic conversation. [26] proposed to represent dual-agent interactions as an optimal control problem, where the actions of the initiating agent induce a cost topology over the space of reactive poses – a space in which the reactive agent plans an optimal pose trajectory. InterFormer [9] uses an interaction transformer with both spatial and temporal attention to generate reactive motions given some initial seed poses of both characters. It incorporates an interaction distance module that integrates softmax-scaled joint distances into the attention matrix. ReMoS [20] proposes a diffusion-based framework and incorporates spatio-temporal cross-attention and hand interaction aware cross-attention. These works primarily focused on generating high-quality short sequences spanning 2 to 5 seconds (around 50 frames), with InterFormer noting an accumulation of errors in longer sequence generation. Ready-to-React [66] introduces a framework that models two-character interactions through independent reactions using an auto-regressive diffusion approach. While Transformer-based approaches produce high-quality results, they face scalability challenges as the number of persons or frames increases, owing to the quadratic complexity of the attention mechanism. In contrast, our method leverages Mamba [21, 13], which alleviates this limitation
@@
State Space Models and Human Motion Synthesis. Despite the well-known superiority of transformer-based architectures in vision tasks  [48, 16, 63, 49, 51, 1, 65, 12, 14, 15, 56, 32, 33, 31, 61, 6, 37, 10, 5, 3, 4], one of their most critical limitations lies in the quadratic complexity of the self-attention mechanism, which results in prohibitive memory consumption and computational overhead when dealing with long sequences or high-resolution visual inputs. This constraint has motivated an increasing body of research aiming to develop alternative sequence modeling paradigms that retain the representational power of transformers while significantly reducing their computational burden.
@@
In this context, state-space sequence models (SSMs), inspired by classical control theory, have recently emerged as a powerful and efficient family of architectures for sequential data modeling [21, 13]. Unlike transformers, which rely on pairwise attention, SSMs leverage parameterized state transitions to capture long-range dependencies in a linear-time manner. Mamba  [21], for instance, introduces a selective SSM formulation that incorporates time-varying parameters into the state-space framework, together with a hardware-aware algorithm design, leading to substantial improvements in both training efficiency and inference scalability.
@@
Thanks to these properties, SSMs have rapidly been adopted across a wide spectrum of visual tasks, ranging from video classification, comprehension, and segmentation  [27, 38, 50, 57], to movie scene detection  [28]. Their applications further extend to cross-modal and generative domains, including text-to-motion generation  [62], image compression  [60], dance video synthesis  [44], and image segmentation  [36, 41, 35, 53], demonstrating their versatility and growing impact in computer vision. Mamba has recently achieved state-of-the-art results in human motion synthesis by generating long, coherent motion sequences [62]. Building on this,  [45] introduced Dyadic Mamba for text-driven dyadic motion generation. Unlike prior approaches, ReactionMamba generates synchronized full-body and hand motions for the reactor using only the actor’s 3D motion, requiring no extra labels or prompts.
@@ III The Proposed Method
Figure 2: Architecture of ReactionMamba. The Mamba encoder encodes the motion of the reaction $\mathbf{Y}$ and maps it to a latent representation $\mathbf{Z}$. The decoder network takes as input the concatenation of the latent vector $\mathbf{Z}$, a projection of the action motion sequence $\mathbf{X}$, and the initial pose $Y_{1}$, and produces the reconstructed reaction motion sequence $\hat{\mathbf{Y}}$.
We aim to generate high-quality reaction motion sequences that are temporally and semantically consistent with a given action motion. The model operates on 3D skeletal pose data, where each pose is a set of joint coordinates over time.
III-A Formulation
Let us consider $P_{t}$ the positions of $k$ distinct joints at time $t$. Consequently, an action sequence $P$ of $T$ frames, can be described as a sequence $P{=}\{P_{1},P_{2},\dots,P_{T}\}$, where $P_{t}{\in}\mathbb{R}^{d}$ and $d{=}3{\times}k$, where $P_{t}{=}[J_{1}(t),\dots,J_{k}(t)],$ with $k$ the number of joints in the skeleton, and $J_{i}(t){=}[x_{i}(t),y_{i}(t),z_{i}(t)]$ the 3D coordinates of joint $i$. The goal is to generate a reaction $\mathbf{Y}{=}\{Y_{1},Y_{2},\dots,Y_{T}\}$, a sequence of skeleton poses, from $\mathbf{X}{=}\{X_{1},X_{2},\dots,X_{T}\}$, a sequence representing the action motion.
The temporal spatial encoder encodes the reaction sequence $\mathbf{Y}$ into a sequence of latent variables $\mathbf{Z}$ through the variational posterior as formulated:
$\mu,\log\sigma^{2}=\mathcal{E}(\mathbf{Y}),\quad\mathbf{Z}=\mu+\epsilon\odot\sigma,\quad\epsilon\sim\mathcal{N}(0,I),$
where $\mathbf{Z}=\{\mathbf{Z}_{1},\dots,\mathbf{Z}_{T}\}$, $\mathbf{Z}_{i}\in\mathbb{R}^{d_{z}}$.
The conditioning module maps the action sequence $\mathbf{X}$ and the initial reaction pose $Y_{1}$ to contextual embeddings $X^{\prime}\in\mathbb{R}^{T\times d_{a}}$ and $Y_{1}^{\prime}\in\mathbb{R}^{T\times d_{p}}$ and then concatenates them with the latent $\mathbf{Z}$, The decoder receives the concatenated sequence $\mathbf{Z}\|X^{\prime}\|Y_{1}^{\prime}]$ and generates the full reaction motion $\hat{\mathbf{Y}}\in\mathbb{R}^{T\times d}$, through the following inverse process:
$\hat{\mathbf{Y}}=\mathcal{D}\left(MLP([\mathbf{Z}\|Y^{\prime}\|X_{1}^{\prime}])\right).$
III-B Preliminaries
SSMs, particularly through the advancements of Structured State Space Sequence Models (S4) and Mamba, have demonstrated exceptional proficiency in modeling long sequences. A Selective Structured State Space Sequence Model (S5) is a neural sequence model derived from a discretized linear time-invariant state space system. Formally, let $u(t)$ denote the input sequence at time $t$, and let $h(t)$ denote the hidden state. The model maps $u(t)$ to an output $v(t)$ using $\mathbf{A}\in\mathbb{R}^{N\times N}$ as the state transition (or evolution) matrix, $\mathbf{B}\in\mathbb{R}^{N\times 1}$ and $\mathbf{C}\in\mathbb{R}^{1\times N}$ as the projection parameters, respectively. The discretized form of the system, incorporating a step size $\Delta$, is expressed as:
$h_{t}$ $=\overline{\mathbf{A}}h_{t-1}+\overline{\mathbf{B}}u_{t},$ (1)
$v_{t}$ $=\mathbf{C}h_{t},$
where $\overline{\mathbf{A}}$ and $\overline{\mathbf{B}}$ are the discretized forms of $\mathbf{A}$ and $\mathbf{B}$.
This formulation enables efficient computation via global convolution, leveraging a structured convolutional kernel $\overline{\mathbf{K}}$ over the full input sequence of length $M$. The kernel is constructed as:
$\overline{\mathbf{K}}$ $=\left(\mathbf{C}\overline{\mathbf{B}},\ \mathbf{C}\overline{\mathbf{A}}\overline{\mathbf{B}},\ \dots,\ \mathbf{C}\overline{\mathbf{A}}^{M-1}\overline{\mathbf{B}}\right),$ (2)
$\mathbf{y}$ $=\mathbf{u}*\overline{\mathbf{K}},$
where * denotes the convolution operation applied along the temporal dimension of the sequence.
@@ III-B Preliminaries
We use Mamba modules as shown in Figure 2 inside both the encoder and decoder to handle temporal dependencies and capture long-range relationships in sequential data.
III-C Model Architecture
The encoder network processes the input reaction motion sequence $\mathbf{Y}$ by first passing it through a linear layer that projects each pose into a hidden dimension $d_{model}$. This projected sequence is then injected into a series of Mamba-based modules Figure 2, directly inspired by the original implementation [42]. Each block consists of a Mamba SSM and a gated multilayer perceptron (MLP) with $d_{intermediate}$ hidden units. The first block receives as input the projected reaction sequence and produces hidden states and residuals that will be used as input of the next block. In each block residual connection fusion applied along with RMS normalization. Once the sequence has passed through all interconnected blocks, a final normalization and a linear projection are used to produce the encoder output of dimensions $(T\times d_{model})$.
The decoder network receives as input the latent vector $\mathbf{Z}$ of dimension $d_{z}$, refined with a projection of the action motion sequence $\mathbf{Y}$ and the initial pose $Y_{1}$ to a space of size $d_{c}$ for each via concatenation. Now this new input is projected to the hidden dimension $d_{model}$. The decoder architecture mirrors that of the encoder, using sequence of Mamba-based modules. Each block again applies a Mamba layer to capture the temporal structure of the input, followed by a gated MLP. Residual connections and RMS normalization are fused at both stages to maintain consistency with the encoder. After processing, a final normalization and linear projection produce the reconstructed reaction motion sequence $\hat{\mathbf{Y}}$.
III-D Loss Functions

### v2

@@ Abstract
We present ReactionMamba, a novel framework for generating long 3D human reaction motions. ReactionMamba integrates a motion VAE for efficient motion encoding with Mamba-based SSMs to decode temporally consistent reactions. This design enables ReactionMamba to generate both short sequences of simple motions and long sequences of complex motions, such as dance and martial arts. We evaluate ReactionMamba on three datasets—NTU120-AS, Lindy Hop, and InterX—and demonstrate competitive performance in terms of realism, diversity, and long-sequence generation compared to previous methods, including InterFormer, ReMoS, Ready-to-React, and Think-Then-React while achieving substantial improvements in inference speed. The code of the paper will be available if the paper is accepted.
@@ I Introduction
In recent years, there has been impressive progress in human motion generation under various user-specified conditions, such as music pieces [54] and text prompts [33]. However, most of these approaches focus on a single subject setting [3, 32] and, as a result, neglect one essential aspect of human motion: rich human-to-human interactions. One important aspect of human interaction is that the action of one person will usually elicit a reaction from another person. Synthesizing human-to-human interactions introduces advanced capabilities for character animation tools and software, with applications across commercial and entertainment media [18], interactive mixed and augmented reality [11], social robotics [50], and avatar interaction [4].
@@
This led recent research to take interest in the task and propose various architectures to generate reaction motions [13, 5, 36]. However, human interaction reactions present a more challenging task due to the motion reactions, which are highly complex and have low periodicity. They require extremely accurate modeling of both spatial and temporal features, without which the generated motion looks unrealistic. While current state-of-the-art methods  [13, 5] can now generate reactive motion for simple and short actions they still struggle with longer sequences. Indeed short sequences often include simple motion such as boxing or kicking that can be more easily modeled due to a direct body contact leading to a clear reaction. Longer sequences however often include more complex interactions, such as dancing, where the relation between the action and the reaction is not always as direct and obvious. Modeling longer interaction sequences requires specific architectures, different from those focused on short sequences, to achieve high quality reaction generation. Precursor works on reaction generation have represented the task as a translation problem which directed them towards the Transformer  [5]. However, this led to low diversity generation and limited performance in the long term. The self-attention mechanism  [43] computes interactions between every pair of tokens in the input sequence. This leads to quadratic time complexity in relation to the sequence length, which becomes a bottleneck for processing long sequences. As a result, it is essential to explore a new architectural paradigm that can handle long-range dependencies while maintaining linear computational complexity to support motion generation tasks. Recent advances have led to renewed interest in State-Space Models (SSMs) [17]. SSMs are a class of mathematical models used for modeling dynamic systems over time, often in fields like control theory, signal processing, and time series analysis. They have several advantages, particularly in their ability to model temporal dependencies for long sequences [17]. The efficiency in managing long sequences primarily due to its implementation of convolutional computations, which efficiently process sequential data [16]. Convolutional operations reduce the complexity associated with handling long-range dependencies by applying local filters to the sequence such as Mamba [15]. In this paper, we propose ReactionMamba (Fig. 2) a new architecture for reaction motion generation that matches the state-of-the-art while increasing inference speed by a hundredfold. Figure LABEL:fig:teaser shows some examples results obtained by our framework on NTU120-AS [47].
@@
  • We present ReactionMamba, a simple yet effective framework that leverages Mamba-based state space models to generate human motion reactions without relying on any auxiliary input or external information.
  • ReactionMamba is a Variational AutoEncoder (VAE) [23] made up of three interconnected modules, an encoder, a conditioning module, and a decoder. The SSM-based encoder encodes the reaction into a sequence of continuous latent variables in an efficient way, and a SSM-based decoder is then designed to capture information hidden in the latent representations of the pose, combined with the action motion sequence and the initial frame of the reaction and to reconstruct the reaction.
  • Unlike conventional deterministic methods [30], our method uses random latent sampling at inference, conditioned by the projected action sequence and the projected initial reaction pose, providing diverse and realistic reaction sequences.
  • The ReactionMamba framework demonstrates strong performance on a range of two-person action–reaction scenarios, including short-duration interaction categories from NTU120-AS [47], longer-duration categorized interactions from InterX [54], and more complex two-person interactions such as Lindy Hop dancing [20], while substantially improving inference speed.
@@ II Related work
Single Subject Human Motion Synthesis. Single subject motion synthesis has seen remarkable progress in recent years. To cite some examples, conditional VAE [32, 7] have been proposed to map textual descriptions to human motion. More recent approaches, such as Editable Dance GEneration (EDGE), [41] and Dancing to Music [24], introduce frameworks for the generation of music-conditioned motion. However, these methods often require additional conditions to effectively model emotional expressiveness. In addition, these approaches focus on the generation of single-human while ReactMamba learns the synchronization between two interacting individuals directly from one person’s motion, without the need for additional prompts or labels.
@@
Reactive Motion Generation. Reactive motion generation focuses on generating human motion in response to external agents. Reactive motion generation is inherently more challenging than generating a single human motion sequence. In [49], a motion graph synthesis technique is proposed that generates gestures and body motions in a dyadic conversation. Huang et al[20] proposed to represent dual-agent interactions as an optimal control problem, where the actions of the initiating agent induce a cost topology over the space of reactive poses – a space in which the reactive agent plans an optimal pose trajectory. Chopin et al[5] proposed InterFormer, an interaction transformer with both spatial and temporal attention to generate reactive motions given some initial seed poses of both characters. It incorporates an interaction distance module that integrates softmax-scaled joint distances into the attention matrix. Ghosh et al[13] introduced Reactive Motion Synthesis (ReMoS), a diffusion-based diffusion framework incorporating spatio-temporal cross-attention and hand-interaction-aware cross-attention. These works primarily focused on generating high-quality short sequences spanning 2 to 5 seconds (around 50 frames), with InterFormer [5] noting an accumulation of errors in longer sequence generation. More recently, Cen et al[1] proposed Ready-to-React, an online reaction policy for two-character interaction generation, where each character independently predicts its next pose in a streaming manner using an auto-regressive model with a diffusion head, enabling long-term interaction generation without requiring future motion information. Tan et al[37] proposed Think-Then-React, an LLM-based framework for online and unconstrained action-to-reaction generation, which introduces an explicit thinking stage to infer action intent and periodically re-thinks during inference to mitigate semantic error accumulation. Although these attention-based approaches produce high-quality results, they face scalability challenges and high memory cost as the number of persons or frames increases, owing to the quadratic complexity of the attention mechanism  [43]. In contrast, our method leverages Mamba [14, 8], which alleviates this limitation.
@@
State Space Models and Human Motion Synthesis. Despite the well-known superiority of transformer-based architectures in vision tasks  [43, 9, 26, 6], one of their most critical limitations lies in the quadratic complexity of the self-attention mechanism, which results in prohibitive memory consumption and computational overhead when dealing with long sequences or high-resolution visual inputs. This constraint has motivated an increasing body of research aiming to develop alternative sequence modeling paradigms that retain the representational power of transformers while significantly reducing their computational burden.
@@
In this context, SSMs, inspired by classical control theory, have recently emerged as a powerful and efficient family of architectures for sequential data modeling [14, 8]. Unlike transformers, which rely on pairwise attention  [43], SSMs leverage parameterized state transitions to capture long-range dependencies in a linear-time manner. Mamba  [14], for instance, introduces a selective SSM formulation that incorporates time-varying parameters into the state-space framework, together with a hardware-aware algorithm design, leading to substantial improvements in both training efficiency and inference scalability.
@@
Thanks to these properties, SSMs have rapidly been adopted across a wide spectrum of visual tasks, ranging from video classification, comprehension, and segmentation  [21, 31, 44, 48], to movie scene detection  [22]. Their applications further extend to cross-modal and generative domains, including text-to-motion generation  [53], image compression  [51], dance video synthesis  [38], and image segmentation  [29, 34, 28, 45], demonstrating their versatility and growing impact in computer vision.
Recently, Zeyu et al[53] proposed Motion Mamba, which achieves state-of-the-art performance in human motion synthesis by generating long, coherent motion sequences. Motion Mamba is a latent-diffusion text-to-motion model that replaces conventional Transformer-based temporal modeling with Mamba (selective state-space models) to effectively process long-range dependencies while maintaining inference efficiency. While Motion Mamba focuses on text-driven single-person motion generation, ReactionMamba addresses reaction synthesis, generating a reactor’s 3D motion in response to an observed actor.
Tanke et al[39] introduced Dyadic Mamba for text-driven dyadic motion generation. Dyadic Mamba employs an SSM/Mamba-based diffusion model to generate text-conditioned, two-person interactions that remain coherent over extended durations. In contrast, ReactionMamba utilizes a VAE-Mamba framework to synthesize a reactor’s 3D motion conditioned on the actor’s movement and the reaction’s initial pose, bypassing the need for text prompts. Furthermore, because Dyadic Mamba relies on iterative multi-step denoising, it entails significant computational costs during inference; ReactionMamba avoids this overhead through its one-step VAE-based formulation.
@@ III The Proposed Method
Figure 2: VAE Architecture of ReactionMamba. The Mamba encoder encodes the motion of the reaction $\mathbf{Y}$ and maps it to a latent representation $\mathbf{z}$. The decoder network takes as input the concatenation of the latent vector $\mathbf{z}$, a projection of the action motion sequence $\mathbf{X}$, and the initial pose $Y_{1}$, and produces the reconstructed reaction motion sequence $\hat{\mathbf{Y}}$.
We aim to generate high-quality reaction motion sequences that are temporally and semantically consistent with a given action motion. The model operates on 3D skeletal pose data, where each pose is a set of joint coordinates over time. The goal is to generate a reaction $\mathbf{Y}{=}\{Y_{1},Y_{2},\dots,Y_{T}\}$, a sequence of skeleton poses, from $\mathbf{X}{=}\{X_{1},X_{2},\dots,X_{T}\}$, a sequence representing the action motion. In this section, we detail the architecture and operational principles of the Reaction Mamba framework, designed for the efficient generation of long-range human reaction motion conditioned on a source action. Initially, we introduce the foundational Mamba model underpinning our approach. We then describe our specialized architecture, which leverages Mamba blocks to improve motion generation efficiency. This framework comprises an encoder for processing input reaction sequences and a decoder that reconstructs the reaction motion.
III-A Preliminaries
Selective Structured State Space Models (SSMs), specifically recent advancements such as S4 [16] and Mamba [15], have demonstrated superior performance in modeling long-range dependencies. Mamba is a selective structured state space model (SSM) that goes beyond the classical discretized linear time-invariant (LTI) state space formulations by introducing input-dependent parameterization, thus enabling information selection over time.
Given an input sequence $\{x_{t}\}_{t=1}^{T}$, Mamba has a latent state $h_{t}\in\mathbb{R}^{N}$ that evolves according to
$h_{t}=\bar{\mathbf{A}}(x_{t})\,h_{t-1}+\bar{\mathbf{B}}(x_{t})\,x_{t},$ (1)
where $\bar{\mathbf{A}}(x_{t})$ and $\bar{\mathbf{B}}(x_{t})$ denote discretized and input-conditioned state transition and input projection matrices, respectively. This selectivity allows the model to dynamically select relevant information over time.
The output is computed as
$y_{t}=\mathbf{C}(x_{t})\,h_{t}+\mathbf{D}\,x_{t},$ (2)
where $\mathbf{C}(x_{t})$ is also an input-dependent output projection and $\mathbf{D}$ represents a learned skip connection that is used for stabilization and optimization purposes.
The discretization of the continuous-time dynamics is done with the help of an input-dependent time-step parameters $\Delta t(x_{t})$, which allow the model to keep or forget information depending on the input. Despite this added complexity, Mamba maintains linear-time complexity $\mathcal{O}(T)$ by using a parallelizable scan operation. This allows the model to be more adapted to the input data and on the same time avoid the quadratic cost of self-attention which allow it to remain effective at modeling long-range dependencies. As illustrated in 2, we integrate Mamba-based modules into both the encoder and decoder to effectively capture long-range temporal relationships within the sequential motion data.
@@ III-A Preliminaries
As illustrated in Fig. 2, we integrate Mamba modules into both the encoder and decoder to effectively capture long-range temporal relationships within the sequential motion data.
III-B Reaction Mamba
The Reaction Mamba VAE $\mathcal{V}$ utilizes a Mamba-based architecture [14, 8] consisting of an encoder $\mathcal{E}$ and a decoder $\mathcal{D}$. We define the mapping and structure of $\mathcal{V}=\{\mathcal{E},\mathcal{D}\}$ as follows.
The encoder network $\mathcal{E}$ processes the input reaction motion sequence $\mathbf{Y}$ by first employing a linear layer to project each pose into a $d_{\text{model}}$-dimensional hidden space. As illustrated in Fig. 2, this projected sequence is processed by a series of Mamba-based modules, following the architecture proposed in [35]. Each block incorporates a Mamba Selective State Space Model (SSM) and a gated multilayer perceptron (MLP) with $d_{\text{intermediate}}$ hidden units.
Within each block, residual connections and Root Mean Square (RMS) normalization are applied to stabilize the hidden state transitions. The initial block operates on the projected reaction sequence, while subsequent blocks ingest the hidden states and residuals from their predecessors. After traversing all interconnected blocks, a final normalization step and a linear projection yield the encoder features of dimension $T\times d_{\text{model}}$. This spatio-temporal encoder maps the reaction sequence $\mathbf{Y}$ into a sequence of latent variables $\mathbf{z}=\{\mathbf{z}_{1},\dots,\mathbf{z}_{T}\}$, where $\mathbf{z}_{i}\in\mathbb{R}^{d_{z}}$. This is formulated through a variational posterior that produces the mean $\mu$ and variance $\sigma$ used for the reparameterization trick during training.
The conditioning module maps the action sequence $\mathbf{X}$ and the initial reaction pose $\mathbf{Y}_{1}$ to contextual embeddings $\mathbf{X}^{\prime}\in\mathbb{R}^{T\times d_{a}}$ and $\mathbf{Y}_{1}^{\prime}\in\mathbb{R}^{T\times d_{p}}$. These embeddings are subsequently concatenated with the latent vector $\mathbf{z}$ to form the joint representation:
$\mathbf{C}=Concat(\mathbf{z},\mathbf{X}^{\prime},\mathbf{Y}_{1}^{\prime}),$ (3)
where $Concat$ denotes the concatenation operation, resulting in a unified input for the subsequent Mamba-based decoding layers.
The decoder network $\mathcal{D}$ reconstructs the reaction motion sequence $\hat{\mathbf{Y}}$ from $\mathbf{C}$ (3). This concatenated representation is subsequently projected into the hidden dimension $d_{\text{model}}$ through a linear layer.
The decoder architecture mirrors the encoder, utilizing a series of Mamba-based modules. Each block comprises a Mamba layer for temporal modeling followed by a gated MLP, with residual connections and RMS normalization applied at each stage to ensure architectural consistency and training stability. Following the final block, a normalization layer and a linear projection are used to map the hidden states back to the motion space. Formally, the decoder generates the reconstructed reaction $\hat{\mathbf{Y}}\in\mathbb{R}^{T\times d}$ through the following process:
$\hat{\mathbf{Y}}=\mathcal{D}\left(\text{MLP}(\mathbf{C})\right),$ (4)
More details about the model are given in Section II of Supplementary material.
III-C Loss Functions